4(sec^2 59°-cot^2 31°)÷3-2÷3sin90°+3tan^2 56°×tan^2 34°=x÷3.find the value of `x`?

Grade:10

1 Answers

Deepak Kumar Shringi
5 years ago
sec259=1+tan259 and tan59=cot31 and tan 56 tan34=1
4/3-2/3+3=11/3 so x=11